OpenStudy (anonymous):

Describe two methods to convert 17/25 to a decimal.

OpenStudy (anonymous):

@HeyitsBeth , @Hollywood_Undead , @joshawad , @marylou004 , @tHe_FiZiCx99

OpenStudy (anonymous):

HELP!!! WILL GIVE FAN AND MEDAL!!

OpenStudy (anonymous):

first, of course, is the division method second is by multiplying the numerator and denomenator with a number such that after multiplying, the denomenator becomes a multiple of 10 e.g. 17/25 * 4/4 = 68/100 = 0.68 done!

OpenStudy (anonymous):

OMG YOUR AMAZING!!!!!!!!!! HOW did you do that that fast???

OpenStudy (anonymous):

my mom and me type really fast

OpenStudy (anonymous):

That's for Sure!!

OpenStudy (anonymous):

but she is way faster than i am so it probly woulda been up faster than that if she did it

OpenStudy (anonymous):

THANKS SOOOOOOOO MUCH!!! Been doing this Math Test For Ike an Hour!! Could U Help With Another??

OpenStudy (anonymous):

hmm maybe??? whats the question??

OpenStudy (anonymous):

Iris is cutting fleece to make scarves for her friends. She has thirteen and one-third feet of fleece to make 5 scarves. How long can she make each scarf? Show your work, and include cancellation in your steps.

OpenStudy (anonymous):

13 1/3 = 40/3 40/3 *1/5 =

OpenStudy (anonymous):

can you do the rest?? i dont wanna just GIVE you the answer

OpenStudy (anonymous):

One sec!!

OpenStudy (anonymous):

3 1/3

OpenStudy (anonymous):

I Got 2 2/3??

OpenStudy (anonymous):

hmmmm... maybe i did something wrong.. ill double check my work one sec

OpenStudy (anonymous):

K

OpenStudy (anonymous):

i keep geting 3 1/3

OpenStudy (anonymous):

i going to use a calculator. Is that OK??

OpenStudy (anonymous):

*I'm